# The Promise of Time-Series Foundation Models for Agricultural Forecasting: Evidence from Marketing Year Average Prices

### v1

@@ Abstract
Forecasting agricultural markets remains a core challenge in business analytics, where nonlinear dynamics, structural breaks, and sparse data have historically limited the gains from increasingly complex econometric and machine learning models. As a result, a long-standing belief in the literature is that simple time-series methods often outperform more advanced alternatives. This paper provides the first systematic evidence that this belief no longer holds in the modern era of time-series foundation models (TSFMs). Using USDA ERS data from 1997-2025, we evaluate 17 forecasting approaches across four model classes, assessing monthly forecasting performance and benchmarking against Market Year Average (MYA) price predictions. This period spans multiple agricultural cycles, major policy changes, and major market disruptions, with substantial cross-commodity price volatility. Focusing on five state-of-the-art TSFMs, we show that zero-shot foundation models (with only historical prices and without any additional covariates) consistently outperform traditional time-series methods, machine learning models, and deep learning architectures trained from scratch. Among them, Time-MoE delivers the largest accuracy gains, improving forecasts by $45\%$ (MAE) overall and by more than $50\%$ for corn and soybeans relative to USDA benchmarks. These results point to a paradigm shift in agricultural forecasting: while earlier generations of advanced models struggled to surpass simple benchmarks, modern pre-trained foundation models achieve substantial and robust improvements, offering a scalable and powerful new framework for highstakes predictive analytics.
@@
Disclaimer: This paper and its contents are not related to Amazon and do not reflect the position of the company and its subsidiaries.
@@ 1 Introduction
Forecasting agricultural commodity prices presents an enduring and complex challenge. Prices are shaped by a confluence of volatile and often unpredictable factors – abrupt weather shocks, evolving political environments, shifting trade policies, and structural changes in global supply chains. These sources of uncertainty generate nonlinear dynamics, regime shifts, and structural breaks that undermine the stability of traditional time series relationships. Prices respond not only to short-term supply and demand fluctuations, but also to long-run structural factors such as technological change and market sentiment, each operating at different temporal scales. Even minor shifts in underlying conditions can generate large and persistent deviations from historical trends. The challenge is compounded by data scarcity. Specialized agricultural markets often lack the high-frequency or long-span datasets needed to train and validate sophisticated forecasting models. Yet accurate price forecasts are crucial for decision-making by farmers, agribusinesses, and policymakers.
@@
Traditional forecasting methods for agricultural prices have employed two main approaches. USDA Economic Research Service (ERS) uses a futures-plus-basis methodology (Tomek, 1997), combining forward-looking market expectations with historical price differentials. Academic and industry forecasters have employed time series models such as ARIMA and VAR (Box and Jenkins, 1970; Sims, 1980), which leverage statistical patterns in historical prices. These approaches, while effective in many contexts, cannot capture complex nonlinear patterns and long-range dependencies in price dynamics due to their linear structure. Recent years have witnessed rapid advances in time series forecasting techniques, moving beyond traditional statistical models toward a new generation of machine learning, deep learning, and more recently, time-series foundational models (TSFM). These approaches promise to capture nonlinear dependencies, high-dimensional feature interactions, and cross-series relationships that conventional econometric models often miss. In particular, time-series foundation models, trained on vast and diverse datasets across domains, claim to offer unprecedented generalization capabilities – adapting to new contexts with minimal or even without any retraining. However, whether these promises translate effectively to the agricultural sector remains largely unexplored. Agricultural data pose distinctive challenges as discussed above. Consequently, it is not yet clear whether these advanced models can deliver meaningful gains in predictive performance over simpler, well-calibrated econometric models in such specialized settings.
@@
Our paper directly addresses this knowledge gap by providing the first comprehensive evaluation of time-series foundation models for agricultural price forecasting. Our analysis proceeds in two complementary stages. First, we conduct a comprehensive comparison of forecasting performance across a wide spectrum of models – ranging from traditional econometric approaches (such as ARIMA and exponential smoothing) to machine learning (e.g., random forests, gradient boosting), deep learning (e.g., LSTM, DeepAR), and the latest time-series foundation models. This stage focuses on monthly price forecasting up to 12 months, enabling us to assess each model’s ability to capture short-term fluctuations, nonlinear dependencies, and evolving market dynamics.
Second, we benchmark the forecasting performance of these models against a policy-relevant metric: the Marketing Year Average (MYA) price. USDA ERS produces official MYA price forecasts that serve as the operational standard for farm program administration, making this a natural benchmark for evaluating alternative forecasting methods. This step moves beyond conventional accuracy measures to evaluate how well models perform when their forecasts are aggregated and applied to a real-world policy context. We focus on the MYA price for two interrelated reasons: its substantive policy relevance and its methodological significance as a challenging yet instructive forecasting target.
@@
From a substantive perspective, MYA prices are among the most policy-relevant indicators in U.S. agriculture. They represent the weighted average of prices received by farmers over the course of a marketing year and serve as the basis for key USDA programs, including the Price Loss Coverage (PLC) and Loan Deficiency Payment (LDP) programs (Zulauf and Schnitkey, 2014; Schnitkey et al., 2019). As such, MYA prices influence farm income, risk management decisions, and government outlays, shaping the broader economic landscape of the agricultural sector. Accurate forecasts of MYA prices are critical for policymakers in setting payment thresholds, for farmers in making planting and storage decisions, and for agribusinesses in managing procurement and contracts. With billions of dollars in farm program payments depending on MYA price forecasts, even modest improvements in forecast accuracy can have substantial economic impacts.
From a methodological perspective, MYA prices offer a stringent test case for evaluating forecasting models. Because they are constructed as weighted averages across months, MYA prices naturally smooth short-term volatility and idiosyncratic shocks. This implies that simpler statistical models (such as AR or exponential smoothing methods) can already deliver reasonable forecasts. The scope for improvement thus appears limited, setting a high bar for more complex algorithms. On the other hand, like other agricultural variables, they remain influenced by nonlinear, multi-scale, and structural forces that may not be fully captured by traditional methods. This dual nature makes MYA prices an ideal setting for examining whether modern machine learning, deep learning, and time-series foundational models can extract subtle patterns and achieve superior predictive accuracy even in aggregated and seemingly stable data environments.
@@
Our evaluation indeed reveals that the conventional wisdom that “simple models forecast best” holds for past deep learning methods – but breaks down decisively with modern pre-trained time-series foundation models, which consistently outperform both traditional and advanced alternatives. Specifically, using data from the USDA ERS on corn, soybeans, wheat, and cotton prices spanning 1997–2025, we evaluate 17 forecasting methods across four categories: traditional time series (ARIMA, ETS, Theta, STL, Prophet, Naive), machine learning (Random Forest, XGBoost), deep learning (LSTM, N-BEATS, TFT, DeepAR), and foundation models (Chronos, Chronos-2, TimesFM 2.5, Time-MoE, Moirai-2). The 28-year sample evaluation period spans multiple agricultural cycles, major policy changes, and pronounced market disruptions – including the 2008 financial crisis, the 2012 drought, and the COVID-19 pandemic – and features substantial cross-commodity variation in price volatility. Through a unified evaluation framework with 1,088 forecasts across 64 train-test splits, we find that foundation models achieve superior performance in monthly price forecasting, with all five ranking in the top five positions. Time-MoE leads with the smallest RMSE and MAE, followed by Chronos and Chronos-2. Deep learning models trained from scratch consistently underperform, ranking 10th–16th despite extensive hyperparameter optimization. The Naive model (rank 6) outperforms all deep learning models, demonstrating that simpler approaches generalize better with limited training data.
@@
For MYA price forecasting, Time-MoE achieves the best overall performance, improving 45.4% over USDA’s operational forecasts in MAE. The gains are particularly large for major row crops: 52.9% on corn and 55.2% on soybeans. Fourteen of seventeen models outperform USDA in this comparison, where forecasts are made before the marketing year starts. Our analysis reveals that smaller mixture-of-experts architectures outperform larger dense transformers, that zero-shot foundation models are competitive with specialized baselines without requiring domain-specific training, and that simpler models generalize better with limited data. These findings have important implications for farm policy design, risk management, and the broader application of foundation models to economics. We offer substantive discussions about our findings in Section 5.
@@
Our work contributes to three strands of literature: agricultural price forecasting, time series foundation models, and the application of machine learning to economic forecasting. [^1] The forecasting literature encompasses far more methods than we can evaluate here, including additional statistical approaches (e.g., TBATS, state-space models), machine learning variants (e.g., LightGBM, CatBoost, support vector regression), deep learning architectures (e.g., WaveNet, Informer, Autoformer, PatchTST), and emerging foundation models (e.g., TimeGPT, Lag-Llama, MOMENT). For comprehensive surveys, see Hyndman and Athanasopoulos (2018) for traditional and machine learning approaches, Makridakis et al. (2018b) for comparative evaluation across method classes, Zhang et al. (2024) for large language models applied to time series, and Kottapalli et al. (2025) for recent developments in time series foundation models. Our goal is not exhaustive comparison but rather to demonstrate that foundation models can be useful for agricultural forecasting and that well-established methods across model classes can outperform USDA’s operational benchmark.
@@
Subsequent research has examined the accuracy of government forecasts. Sanders and Manfredo (2008) documented significant market reactions to USDA reports for wheat, soybeans, and hogs, while Isengildina-Massa et al. (2011) compared USDA forecasts with private sector predictions across corn, soybeans, wheat, and cotton, finding that USDA maintains competitive accuracy despite using simpler methodologies – a result that motivates our focus on whether modern methods can improve upon this established benchmark.
@@ 2.1 Data Sources and Coverage
Our analysis uses monthly price data from the USDA Economic Research Service (ERS) database for Season-Average Price Forecasts, covering major U.S. agricultural commodities that collectively represent over 200 million planted acres annually. Table 1 provides a comprehensive overview of the data structure, variable definitions, and sources.
@@
Variable Definition Unit Frequency Source Panel A: Primary Variables Monthly Price Price received by farmers \$/bu (corn, soy, wheat) Monthly USDA NASS for commodity sales cents/lb (cotton) Agricultural Prices Panel B: Benchmark Forecasts USDA MYA Price Forecast Official MYA price forecast \$/bu or cents/lb Monthly USDA ERS (futures + basis method) Season-Avg Forecasts Panel C: Auxiliary Variables Futures Price Nearby futures contract \$/bu or cents/lb Daily LSEG Data & Analytics settlement price (formerly Refinitiv) Basis Cash price - futures price \$/bu or cents/lb Monthly USDA ERS (5-year or 7-year average) (calculated) Marketing % Proportion of annual production marketed in each month Percentage Monthly USDA NASS Agricultural Prices
Notes: This table summarizes all variables used in the analysis. Panel A shows the primary variable (monthly prices received by farmers) used to generate forecast. Panel B shows the benchmark forecast used for comparison. Panel C shows auxiliary variables used only by baseline methods (USDA futures-plus-basis approach). All data are publicly available from USDA sources. Marketing percentages are fixed weights (5-year average for corn/soybeans/wheat, 7-year Olympic average for cotton) used to aggregate monthly prices to MYA. See Appendix A for complete data access details and download links.
We use the latest dataset as of 12/10/2025, which includes four major U.S. agricultural commodities: corn (September 1997 – August 2025), soybeans (September 1997 – August 2025), wheat (June 1997 – August 2025), and cotton (August 1997 – August 2025). Each commodity follows its specific marketing year calendar defined by USDA: corn and soybeans (September–August), wheat (June–May), and cotton (August–July). These marketing years align with harvest timing and traditional marketing patterns for each crop.
@@
2.2 Marketing Year Average Prices
The Marketing Year Average (MYA) price is the primary policy-relevant metric for U.S. agricultural programs. MYA represents the weighted average price farmers receive over the 12-month marketing year, calculated as:
MYA $=\sum_{t=1}^{12}P_{t}\times w_{t},$ (1)
where $P_{t}$ is the monthly price received by farmers in month $t$ of the marketing year, and $w_{t}$ is the monthly marketing percentage – the proportion of annual production marketed in month $t$, with $\sum_{t=1}^{12}w_{t}=1$.
Marketing percentages reflect the typical temporal pattern of commodity sales throughout the year. For example, corn marketing percentages are highest immediately post-harvest (September–November: 45% of annual sales) and decline through the marketing year as on-farm storage is depleted. Following USDA convention, each month’s marketing percentage is calculated as the 5-year average of that same month’s historical percentage for corn, soybeans, and wheat, and a 7-year Olympic average (dropping the highest and lowest values) for cotton.
MYA prices determine Farm Bill program payments worth billions of dollars annually through multiple mechanisms. Price Loss Coverage (PLC) is triggered when MYA falls below the statutory reference price, providing counter-cyclical support to farmers during periods of low prices. Agricultural Risk Coverage (ARC) bases payments on MYA relative to a 5-year Olympic average (dropping the highest and lowest values before averaging), protecting against revenue declines. Additionally, MYA prices inform premium rates and indemnity calculations for crop insurance revenue protection policies.
Accurate MYA forecasts enable farmers to make informed planting decisions, help policymakers estimate program costs, and allow financial institutions to assess agricultural credit risk. This makes MYA forecasting accuracy the primary evaluation criterion for our study.
2.3 USDA Official MYA Price Forecast
The USDA ERS has produced MYA price forecasts and updated them every month since 2003 using a futures-plus-basis methodology. For months within the marketing year where actual prices are not yet available, forecasts are generated as:
$\text{Forecast}_{t}$ $=\text{Futures}_{t}+\overline{\text{Basis}}_{t},$ (2)

### v2

# The Promise of Time-Series Foundation Models for Agricultural Forecasting: Evidence from Commodity Prices

@@ Abstract
Forecasting agricultural markets remains challenging due to nonlinear dynamics, structural breaks, and sparse data. A long-standing belief holds that simple time-series methods outperform more advanced alternatives. This paper provides the first systematic evidence that this belief no longer holds with modern time-series foundation models (TSFMs). Using USDA ERS monthly commodity price data from 1997-2025, we evaluate 17 forecasting approaches across four model classes, including traditional time-series, machine learning, deep learning, and five state-of-the-art TSFMs (Chronos, Chronos-2, TimesFM 2.5, Time-MoE, Moirai-2), and construct annual marketing year price predictions to compare with USDA’s futures-based season-average price (SAP) forecasts. We show that zero-shot foundation models consistently outperform traditional time-series methods, machine learning, and deep learning architectures trained from scratch in both monthly and annual forecasting. Furthermore, foundation models remarkably outperform USDA’s futures-based forecasts on three of four major commodities despite USDA’s information advantage from forward-looking futures markets. Time-MoE delivers the largest accuracy gains, achieving 54.9% improvement on wheat and 18.5% improvement on corn relative to USDA ERS benchmarks on recent data (2017-2024 excluding COVID). These results point to a paradigm shift in agricultural forecasting.
@@ 1 Introduction
Forecasting agricultural commodity prices presents an enduring and complex challenge. Prices are shaped by a confluence of volatile and often unpredictable factors – abrupt weather shocks, evolving political environments, shifting trade policies, and structural changes in global supply chains. These sources of uncertainty generate nonlinear dynamics, regime shifts, and structural breaks that undermine the stability of traditional time series relationships. Prices respond not only to short-term supply and demand fluctuations, but also to long-run structural factors such as technological change and market sentiment, each operating at different temporal scales. Even minor shifts in underlying conditions can generate large and persistent deviations from historical trends. The challenge is compounded by data scarcity. Specialized agricultural markets often lack the high-frequency or long-span datasets needed to train and validate sophisticated forecasting models. Yet, accurate price forecasts are crucial for decision-making by farmers, agribusinesses, and policymakers.
@@
Traditional forecasting methods for agricultural prices have employed two main approaches. USDA Economic Research Service (ERS) uses a futures-plus-basis methodology (Hoffman, 2005; Hoffman et al., 2007, 2015; Hoffman and Meyer, 2018), combining forward-looking market expectations with historical price differentials. Academic and industry forecasters have employed time series models such as ARIMA and VAR (Box and Jenkins, 1970; Sims, 1980), which leverage statistical patterns in historical prices. These approaches, while effective in many contexts, cannot capture complex nonlinear patterns and long-range dependencies in price dynamics due to their linear structure. Recent years have witnessed rapid advances in time series forecasting techniques, moving beyond traditional statistical models toward a new generation of machine learning, deep learning, and more recently, time-series foundational models (TSFM). These approaches promise to capture nonlinear dependencies, high-dimensional feature interactions, and cross-series relationships that conventional econometric models often miss. In particular, time-series foundation models, trained on vast and diverse datasets across domains, claim to offer unprecedented generalization capabilities – adapting to new contexts with minimal or even without any retraining. However, whether these promises translate effectively to the agricultural sector remains largely unexplored. Agricultural data pose distinctive challenges as discussed above. Consequently, it is not yet clear whether these advanced models can deliver meaningful gains in predictive performance over simpler, well-calibrated econometric models in such specialized settings.
Our paper directly addresses this knowledge gap by providing the first comprehensive evaluation of time-series foundation models for agricultural price forecasting. Prior evaluations of agricultural forecasting methods suffer from heterogeneity in methodologies, commodities, and sample periods, making systematic comparison difficult (Isengildina-Massa et al., 2024). We address this limitation through a unified evaluation framework: 17 models across four commodities using identical data, forecast horizons, and metrics, enabling direct performance comparisons. From a methodological perspective, our study contributes to understanding when and why foundation models succeed in specialized domains. Agricultural prices present an ideal test case: limited training data (200-400 monthly observations), strong domain-specific patterns (biological production cycles, policy-driven structural breaks), and an established operational benchmark (USDA ERS forecasts). Our findings that zero-shot foundation models outperform domain-specialized methods have implications for forecasting in other data-scarce economic domains.
@@
Our analysis proceeds in two complementary stages. First, we conduct a comprehensive comparison of forecasting performance across a wide spectrum of models, ranging from traditional econometric approaches (such as ARIMA and exponential smoothing) to machine learning (e.g., random forests, gradient boosting), deep learning (e.g., LSTM, DeepAR), and the latest time-series foundation models. This stage focuses on monthly price forecasting up to 12 months, enabling us to assess each model’s ability to capture short-term fluctuations, nonlinear dependencies, and evolving market dynamics.
Second, we benchmark the forecasting performance of these models against a policy-relevant metric: the Marketing Year Average (MYA) price, calculated as the season-average of monthly prices weighted by marketing percentages (hence USDA’s term ”Season-Average Price” or SAP). USDA ERS produces SAP forecasts using a transparent, model-based methodology (futures-plus-basis) that serves as a key input to official WASDE reports. We use ERS SAP forecasts as our benchmark because they are publicly available and replicable, unlike WASDE SAP forecasts which incorporate confidential expert judgment and analyst research. Throughout this paper, we use ”MYA” to refer to the price metric itself and our model forecasts of it, while ”SAP” refers specifically to USDA ERS forecasts of the same metric. Both terms describe the identical concept: the marketing-year weighted average of monthly prices.
@@
We focus on MYA prices for two interrelated reasons: substantive policy relevance and methodological significance as a challenging forecasting target. From a substantive perspective, MYA prices are among the most policy-relevant indicators in U.S. agriculture. They represent the weighted average of prices received by farmers over the marketing year. ERS SAP forecasts serve as key inputs to official WASDE SAP forecasts, which in turn are used to calculate Price Loss Coverage (PLC) and Agricultural Risk Coverage (ARC) payment rates for program commodities (Hoffman and Meyer, 2018). WASDE reports guide production, investment, and trading decisions by farmers, agribusinesses, policymakers, and other stakeholders. With billions of dollars in farm program payments depending on these price forecasts, even modest improvements in forecast accuracy can have substantial economic impacts.
From a methodological perspective, MYA prices offer a stringent test case for evaluating forecasting models. The USDA ERS benchmark incorporates futures prices from actively traded markets where participants have strong financial incentives to form accurate expectations (Fama, 1970). A substantial body of research demonstrates the superiority of futures-based approaches over statistical time-series methods across diverse commodities and horizons (Kastens and Schroeder, 1998; Manfredo and Sanders, 2004; Roache and Reichsfeld, 2011; Hoffman et al., 2015; Etienne et al., 2019; Figuerola-Ferretti et al., 2021). This performance advantage arises because futures markets aggregate dispersed information about anticipated supply, demand, weather, and policy conditions (Adjemian and others, 2020), making them difficult to outperform with backward-looking statistical models. The scope for improvement thus appears limited, setting a high bar for alternative forecasting methods. This makes MYA prices an ideal setting for examining whether modern machine learning, deep learning, and time-series foundation models can achieve superior predictive accuracy even when competing against informationally efficient market-based benchmarks.
@@
Our evaluation indeed reveals that the conventional wisdom that “simple models forecast best” holds for past deep learning methods – but breaks down decisively with modern pre-trained time-series foundation models, which consistently outperform both traditional and advanced alternatives. Specifically, using data from the USDA ERS on corn, soybeans, wheat, and cotton prices spanning 1997–2025, we evaluate 17 forecasting methods across four categories: traditional time series (ARIMA, ETS, STL, Prophet, Naive), machine learning (Random Forest, XGBoost), deep learning (LSTM, N-BEATS, TFT, DeepAR), and foundation models (Chronos, Chronos-2, TimesFM 2.5, Time-MoE, Moirai-2). The 28-year sample evaluation period spans multiple agricultural cycles, major policy changes, and pronounced market disruptions – including the 2008 financial crisis, the 2012 drought, and the COVID-19 pandemic – and features substantial cross-commodity variation in price volatility. Through a unified evaluation framework with 1,088 forecasts across 64 train-test splits, we find that foundation models achieve superior performance in monthly price forecasting, with all five ranking in the top five positions. Time-MoE leads with the smallest RMSE and MAE, followed by Chronos and Chronos-2. Deep learning models trained from scratch consistently underperform, ranking 10th–16th despite extensive hyperparameter optimization. The Naive model (rank 6) outperforms all deep learning models, demonstrating that simpler approaches generalize better with limited training data.
@@
For MYA forecasting, foundation models demonstrate strong performance across major commodities despite using only historical prices while USDA incorporates forward-looking futures market information. On wheat, 13 of 17 models outperform USDA’s SAP prediction, led by Time-MoE with 54.9% improvement in mean absolute error (MAE). On corn and soybeans, foundation models achieve selective but meaningful advantages, with Time-MoE improving 18.5% on corn and TimesFM improving 6.9% on soybeans. Cotton remains challenging, with USDA maintaining superior accuracy. Our analysis reveals that smaller mixture-of-experts architectures outperform larger dense transformers, that zero-shot foundation models are competitive with specialized baselines without requiring domain-specific training, and that simpler models generalize better with limited data. These findings have important implications for farm policy design, risk management, and the broader application of foundation models to economics.
@@
Our work contributes to three strands of literature: agricultural price forecasting, time series foundation models, and the application of machine learning to economic forecasting. [^1] The forecasting literature encompasses far more methods than we can evaluate here, including additional statistical approaches (e.g., TBATS, state-space models), machine learning variants (e.g., LightGBM, CatBoost, support vector regression), deep learning architectures (e.g., WaveNet, Informer, Autoformer, PatchTST), and emerging foundation models (e.g., TimeGPT, Lag-Llama, MOMENT). For comprehensive surveys, see Hyndman and Athanasopoulos (2018) for traditional and machine learning approaches, Makridakis et al. (2018b) for comparative evaluation across method classes, Zhang et al. (2024) for large language models applied to time series, and Kottapalli et al. (2025) for recent developments in time series foundation models. Our goal is not exhaustive comparison but rather to demonstrate that foundation models can be useful for agricultural forecasting and that well-established methods across model classes can be competitive with USDA’s ERS SAP forecasts.
@@
Subsequent research has examined the accuracy of government forecasts. Sanders and Manfredo (2008) documented significant market reactions to USDA reports for wheat, soybeans, and hogs, while Isengildina-Massa et al. (2011) compared USDA forecasts with private sector predictions across corn, soybeans, wheat, and cotton, finding that USDA maintains competitive accuracy despite using simpler methodologies. Systematic reviews (Isengildina-Massa et al., 2024) document that USDA forecasts consistently outperform naive and time-series alternatives and achieve accuracy comparable to market-based forecasts, establishing them as a demanding benchmark for methodological innovation. This result motivates our focus on whether modern methods can improve upon this established benchmark.
The USDA ERS futures-plus-basis methodology has evolved through systematic research by Hoffman and colleagues. Hoffman (2005) developed the initial futures-based model for corn, demonstrating that adding 5-year average basis to futures prices produces reliable forecasts. Hoffman et al. (2007) extended this approach to soybeans and wheat, finding that 5-year moving averages effectively address structural changes while maintaining forecast precision. Hoffman and Meyer (2018) adapted the methodology for cotton using 7-year Olympic averages (excluding maximum and minimum values) to improve basis forecast stability. This methodological development reflects broader empirical evidence that futures-based forecasts consistently outperform time-series alternatives across agricultural commodities (Kastens and Schroeder, 1998; Manfredo and Sanders, 2004; Roache and Reichsfeld, 2011; Hoffman et al., 2015; Etienne et al., 2019; Figuerola-Ferretti et al., 2021). The performance advantage arises because futures markets aggregate dispersed information about anticipated supply, demand, weather, and policy conditions (Adjemian and others, 2020), making them difficult to outperform with backward-looking statistical models.
@@ 2.1 Data Sources and Coverage
Our analysis uses monthly price data from the USDA Economic Research Service (ERS) database for SAP Forecasts, covering major U.S. agricultural commodities that collectively represent over 200 million planted acres annually. Table 1 provides an overview of the data structure, variable definitions, and sources.
@@
Variable Definition Unit Frequency Source Panel A: Primary Variables Monthly Price Price received by farmers \$/bu (corn, soy, wheat) Monthly USDA NASS for commodity sales cents/lb (cotton) Agricultural Prices Panel B: Benchmark Forecasts USDA SAP Forecast ERS SAP forecast \$/bu or cents/lb Monthly updates USDA ERS (futures + basis method) (annual target) Season-Avg Forecasts Panel C: Auxiliary Variables Futures Price Nearby futures contract \$/bu or cents/lb Daily LSEG Data & Analytics settlement price (formerly Refinitiv) Basis Cash price - futures price \$/bu or cents/lb Monthly USDA ERS (5-year or 7-year average) (calculated) Marketing % Proportion of annual production marketed in each month Percentage Monthly USDA NASS Agricultural Prices
Notes: This table summarizes all variables used in the analysis. Panel A shows the primary variable (monthly prices received by farmers) used to generate forecast. Panel B shows the benchmark forecast used for comparison. ERS SAP forecasts are generated weekly using data available each Thursday; files are updated monthly and posted 1 business day after WASDE report release. Panel C shows auxiliary variables used only by baseline methods (USDA futures-plus-basis approach). Marketing percentages are fixed weights (5-year average for corn/soybeans/wheat, 7-year Olympic average for cotton) used to aggregate monthly prices to MYA. All data are publicly available from USDA sources. See Appendix A for complete data access details and download links.
We use the latest dataset as of 12/10/2025, which includes four major U.S. agricultural commodities: corn (September 1997 – August 2025), soybeans (September 1997 – August 2025), wheat (June 1997 – August 2025), and cotton (August 1997 – August 2025). Each commodity follows its specific marketing year calendar [^2] Marketing years are 12-month periods aligned with each commodity’s harvest and marketing cycle, used by USDA for reporting production, sales, and price statistics. defined by USDA: corn and soybeans (September–August), wheat (June–May), and cotton (August–July). Throughout this paper, we denote marketing years by their starting year. For example, MY 2023 for corn refers to the period September 2023 through August 2024. These marketing years align with harvest timing and traditional marketing patterns for each crop.
@@
2.2 USDA ERS Season-Average Price Forecasts
The USDA ERS produces futures-based season-average price (SAP) forecasts that serve as key inputs to the official WASDE consensus forecasts (Poghosyan et al., 2025). WASDE reports combine analyst expertise, econometric models, and market intelligence to generate official price projections. To prevent market manipulation, WASDE preparation follows strict confidentiality protocols, with no information released until the scheduled publication time. In contrast, the ERS SAP forecasts are publicly documented and methodologically transparent, making them suitable as a replicable benchmark. The USDA ERS has produced SAP forecasts since 2003 using a futures-plus-basis methodology in two steps (Hoffman, 2005; Hoffman et al., 2007, 2015; Hoffman and Meyer, 2018).
The ERS methodology first generates monthly price forecasts using futures prices and historical basis patterns. The forecast for month $k$ in year $t$ is generated as:
$\widehat{P}_{k,t}$ $=F_{h(k),t,i}+\bar{B}_{k,t},$ (1)